y=-5/7x-10/9 Find slope and y intercept of line write the fractional answer in lowest terms.
Found 2 solutions by fractalier, mananth:
Answer by fractalier(6550)

Using the slope-intercept form of a line:
y = mx + b
you can see that for
y=-5/7x-10/9
m = slope = -5/7
and
b = y-intercept = -10/9

Answer by mananth(16946)

y=-5/7x-10/9
...
slope = (-5/7)
..
y intercept = (-10/9)
...
m.ananth@hotmail.ca